# Association of the IceCube neutrinos with CAZ blazar light curves

### v1

@@ 4 Results and discussion
| Blz | Type | # | Var # | BB95 | Prominent BBHOP | BB95 as BBHOP peak | |
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | |
(all) | 3225 | 1170 | 8 2.26 (=) | 55 6.19 (=) | 3 1.30 (=) | |
| FSRQ | 1773 | 716 | 3 0.81 (=) | 32 3.01 (=) | 0 0.00 (=) | ||
| BLL | 296 | 212 | 3 1.31 (=) | 18 2.91 (=) | 2 1.30 (=*) | ||
(all) | 3814 | 1917 | 14 3.44 (=) | 85 11.15 (=) | 4 1.37 (=) | |
| FSRQ | 792 | 504 | 1 0.11 (=) | 21 2.18 (=) | 0 0.00 (=) | ||
@@
We show the results of the spatio-temporal correlation analysis between the CAZ optical light curves and the IceCube neutrinos in Table 2. We list all individual spatio-temporal associations with (i.e., greater than the median weight) in Table 3. The full list of the spatial and spatio-temporal blazar-neutrino associations is given as an electronic table. The associations are visualized in an all-sky map in Fig. 1.
As described in Sect. 2.2, the test is performed on two samples of blazars: RFC† and 4LAC. Since FSRQs and BLLs are thought to be physically distinct (see Sect. 1), several studies investigated their correlation with neutrinos separately (e.g., Aartsen et al. 2017a). As such, in this paper we additionally test the spatio-temporal correlation of neutrinos with FSRQs and BLLs separately. This results in a total of six different blazar subsamples being tested, hence, six rows in Table 2. As described in Sects. 2.2.4, 2.2.5, and 2.2.6, the temporal association is tested via three metrics: (1) BB95 periods; (2) prominent BBHOP flares; and (3) BB95 periods forming the peak of a prominent BBHOP flare. Thus, altogether we have 18 -values.
Overall, we find that the total number of high-weight blazar-neutrino spatio-temporal associations is low. As a result, blazars and neutrinos exhibit a weaker than 2 (pre-trial) spatio-temporal correlation in 14 out of the 18 tested scenarios. Of the four scenarios crossing the 2 threshold, three are weaker than 3, while one reaches 3.48 (pre-trial). Crucially, the strongest correlations are driven by only two high-weight individual associations, namely: (1) the IBL CAZJ0211+1051 (4FGL J0211.2+1051) with the neutrino IC131014A (=); and (2) the IBL CAZJ0509+0541 (TXS 0506+056) with the well-known IC170922A (=).
@@ 4.1 Effect of seasonal gaps
The low number of spatio-temporal associations could be due to a major data limitation present in the CAZ light curves: frequent and long seasonal gaps. Of the 605 spatial associations between neutrinos and variable blazars, in 333 (55.0%) the neutrino arrives in a part of the light curve which is without data. Either the neutrino arrives outside of the total range of the light curve (in 79 cases) or within a seasonal gap (in 172 cases in the CRTS portion, and 82 in the rest of the light curve).
While it is not known which of these spatially neutrino-associated blazars would have been flaring at the time of the neutrino arrival, we can estimate the fraction. For the 272 (45.0% of 605) spatial associations where we have data, in 110 (39.7% of 272) the neutrino coincides with a BB95 period or a prominent BBHOP flare. Unsurprisingly, this coincidence rate is similar to the combined average duty cycle of prominent BBHOP flares (34%; see Fig. 12 of K25-Cat) and that of BB95 (5%; see Fig. 14 of of K25-Cat). If we assume such a flare coincidence rate for all variable light curves globally, we estimate that around potential spatio-temporal associations are missed because of the seasonal gaps. This is a notable caveat for our spatio-temporal analysis because it prevents us from identifying a potential temporal association in around of the spatial associations.
@@ 4.2 BB95 periods versus prominent BBHOP flares
As seen from Table 2, using BB95 periods as a proxy for temporal associations leads to substantially stronger, albeit fewer, correlations than using prominent BBHOP flares. As described in Sect. 2.2.4, BB95 periods are exclusive to the top 5% flux densities and can be as short as a few days. On the other hand, as described in Sect. 2.2.5, BBHOP flares trace a major outburst, whose peak reaches the 75th percentile flux density, from its inception to its termination regardless of the flux density at the start and end. BBHOP flares can be as long as hundreds of days.
@@
Interestingly, the strongest correlation strengths are found upon minimizing the effect of the noise-driven variations. This is done by temporally associating neutrinos with BB95 periods which form the peak of a prominent BBHOP flare (see Sect. 2.2.6). Since this metric requires a relatively high cadence to be identified, they are generally rare in the CAZ light curves. Still, two of the five blazar-neutrino associations emerging from this metric are with the two most high-weight spatio-temporal associations, which greatly enhances the correlation strengths obtained via this temporal metric. This reinforces the above interpretation that neutrinos preferably arrive at the peak of the largest outburst periods from blazars.
@@ 4.3 FSRQs versus BLLs
When testing blazar-neutrino correlations with blazars divided into FSRQs and BLLs, Table 2 shows that BLLs are more strongly spatio-temporally correlated with the updated IceCat1+ neutrinos than FSRQs. In both RFC† and 4LAC samples, and via all the temporal metrics, the correlation strength substantially increases when going from all blazars to BLLs, while a decrease is seen when going from all blazars to FSRQs. This generally contrasts the results that suggest FSRQs are more strongly correlated with neutrinos (e.g., Moretti & Caccianiga 2025).
Naively, our result would suggest that BLLs are better neutrino emitters than FSRQs. However, this generalization should be considered cautiously, because the correlation is driven by only two individual high-weight associations, namely: (1) the IBL CAZJ0211+1051 with IC131014A; and (2) the IBL CAZJ0509+0541 with IC170922A. First, it is possible that we are missing other potential spatio-temporal associations due to the presence of long and frequent seasonal gaps in the CAZ light curves (see Sect. 4.1). With more data, it is possible that this apparent BLL exclusivity disappears. Second, these correlations could arise because of some specific properties present in these individual sources, possibly unique to them rather than their entire class. Third, this generalization is strongly dependent on the reliability of the source classifications. If these blazars are misidentified, they can falsely favor the misidentified class as the more likely neutrino emitter because of their high weight. On this regard, the BLL-classification of some neutrino-emitting candidate blazars have been questioned in the past (e.g., that of TXS 0506+056, Padovani et al. 2019).
@@ 4.4 Fraction of cosmic neutrinos emitted by blazars
In K25-Sim we found that even if only 20% of neutrinos were emitted by blazars at the time of a major flare, our most optimal spatio-temporal test setup (counted and top-hat weighted; used in this study) should confidently detect the correlation. The simulations had 4000 blazars with a 30% flaring rate, comparable to the RFC† and 4LAC tests here. Of the 1000 simulations we ran, in 87.9% the blazar-neutrino spatio-temporal correlation was detected at least at the 4 level, and in 97.8% at the 3 level. The mean of the 1000 -values was 0.0004. To compare the observed significances to the simulation ones, we focus on -values obtained via the BB95 temporal metric. This is because the BB95 temporal metric uses the extreme tail of the flux density distribution to identify flares, similar to how flares were defined in the simulations. Thus, the observed -value for all RFC† blazars is 0.1246 (i.e., 1.54) and for all 4LAC blazars 0.2038 (i.e., 1.27). Clearly, fewer than 20% of the cosmic neutrinos were emitted during major optical flares of blazars.
@@
There are two differences between the tests here and those of K25-Sim: (1) the effect of seasonal gaps was not simulated in K25-Sim, whereas here the light curves have a seasonal gap rate of 55% leading to the omission of 22% of spatial associations as potential spatio-temporal ones (see Sect. 4.1); and (2) in K25-Sim the original IceCat1+ neutrino sample (283 events) was used, but here we use the updated version (356 events; larger by 26%). While the effect of (1) may be critical and needs to be accounted for, we checked that the effect of (2) is minimal by repeating the simulations of K25-Sim using the updated IceCat1+.
@@
To reliably estimate the fraction of astrophysical neutrinos emitted by blazars during major optical flares, here we run similar simulations to those of K25-Sim while taking into account the seasonal gaps. We generate 4000 random blazars with a 30% chance of being variable and a few tens of variable blazars simulated to be spatio-temporally associated with neutrinos. The exact number of simulated blazar-neutrino associations depends on the signalness of the neutrinos and varies from one simulation run to another. We then apply the spatio-temporal correlation analysis to the updated IceCat1+ neutrinos and these simulated blazars. We focus on the spatio-temporally counted, top-hat-weighted scenario, as used in the real tests of this study. We simulate the effect of the 55% light curve gap fraction by randomly excluding 50% of spatio-temporal associations from the -value calculations. Next, we incrementally increase the fraction of blazar-flare-emitted cosmic neutrinos (i.e., those simulated to have been spatio-temporally associated with blazars) from 0% to 20%. These simulations are repeated 500 times. We note that the Monte Carlo number for each of the 500 repetitions is , meaning that the smallest obtainable -value is 0.0001 (i.e., 3.9; similar to K25-Sim).
@@
In Fig. 2, we plot the correlation strength of the simulations in Gaussian against the fraction of cosmic neutrinos emitted during major blazar flares in the optical band. Notably, when this fraction is 20%, the typical correlation strengths are 4, similar to those from K25-Sim. This suggests that excluding even as much as 50% of the spatio-temporal associations from the -value calculations does not seem to affect the -values by much. Therefore, while the presence of seasonal gaps can substantially reduce the total number of spatio-temporal associations, it only modestly weakens the overall spatio-temporal correlation strengths. Nevertheless, we caution that this conclusion may only be valid within the assumed 0%–20% fraction of cosmic neutrinos from blazar flares.
@@
In Fig. 2, where the gap fraction is taken as 50%, we additionally show the observed correlation strengths for all RFC† and 4LAC blazars via the most comparable temporal metric (i.e., BB95). These are 1.54 and 1.27, respectively. As evident, 1.54 spatio-temporal correlation strengths are most often obtained when 2–8% of cosmic neutrinos are emitted during major blazar flares. Based on the 4LAC tests, this fraction would be even smaller. This fraction would also be smaller if the light curve gap fraction were smaller than 50%. For example, for a light curve gap fraction of 20%, the observed fraction of cosmic neutrinos from blazar flares would be 5%. Thus, we can confidently conclude that the fraction of cosmic neutrinos emitted during major optical flares of blazars is 8%. Remarkably, this is well in agreement with the previous estimates for the fraction of neutrinos coming from blazars (e.g., Aartsen et al. 2017b; Huber 2019; Oikonomou 2022). Interestingly, even among other neutrino-emitting source candidates, like Seyfert-II AGN, neutrino emission appears to be dominated by only a few sources rather than the whole population (e.g., Saurenhaus et al. 2025).
@@
If indeed 8% astrophysical neutrinos come during blazar optical flares, it is then expected that only 12 neutrinos from the updated IceCat1+ sample were emitted during bright optical flares. Due to the previously discussed light curve limitations, we unfortunately cannot confidently identify most of their potential emitters, except for CAZJ0211+1051 and CAZJ0509+0541.
4.5 Individual spatio-temporal blazar-neutrino associations
@@
Here we individually investigate a few interesting spatio-temporal blazar-neutrino associations. We begin with all four associations arising from the BB95 metric with neutrinos that have . This threshold selects the most reliably reconstructed neutrino events and is similar to a common quality threshold used in the literature, namely: and . In the updated IceCat1+ sample there are 79 (22% of 356) neutrino events that meet , of which 64 (81%) are classified as “gold” events by the IceCube collaboration. Their minimum, median, and maximum (and ) are 0.502, 0.663, and 0.997, respectively. Likewise, their minimum, median, and maximum are 3.3, 5.2, and 12.3 , respectively.
@@
Two of the four BB95 associations with (namely, the IBL CAZJ0211+1051 with IC131014A, and the IBL CAZJ0509+0541 with IC170922A) also fulfill the BB95 as the peak of a BBHOP temporal metric. These two associations are the most prominent in our study. In Appendix A, we also explore three additional associations that arise from this metric regardless of their low . The complete list of spatial and spatio-temporal associations is given in an electronic table.
@@
Dedicated studies on CAZJ0211+1051 are needed to confirm if it really is a neutrino emitter. Recent studies have suggested a link between neutrino emission and changes in the multiwavelength polarization degree and angle of blazars (directly, e.g., Paraschos et al. 2025; and indirectly, e.g., Novikova et al. 2023; Blinov & Novikova 2025). Additionally, polarization measurements can reliably trace shocked regions in the jet (e.g., Liodakis et al. 2022b; Kouch et al. 2024b) that could accelerate protons (a key ingredient of unraveling the blazar-neutrino mystery; e.g., Stathopoulos & Petropoulou 2025). Interestingly, CAZJ0211+1051 is one of the prime targets of multiwavelength photo-polarimetric studies aimed at distinguishing between leptonic and hadronic emissions in blazar jets (e.g., Peirson et al. 2022; Kouch et al. 2025c; Agudo et al. 2025) and is known to exhibit extreme, rapid changes in its polarization properties (e.g., Liodakis et al. 2024). Thus, by combining such photo-polarimetric studies with dedicated polarization monitoring, future studies could confirm the suspected neutrino-emitting nature of CAZJ0211+1051.
The next spatio-temporal association we focus on is that of CAZJ0207+0950 (4FGL J0207.9+0953; a blazar candidate) with IC131014A, which is, as discussed above, also spatio-temporally associated with the neighboring blazar CAZJ0211+1051. As seen in Fig. 4, the neutrino arrived 20 d after the peak of a major flare. Notably, while this meets both the BB95 and BBHOP metrics separately, it does not meet the BB95 at the peak of a BBHOP metrics since the neutrino did not arrive exactly at the peak of the BBHOP flare. Due to the statistical nature of the neutrino data, it is not possible to definitively determine from which (if any) of the two sources this neutrino emerged. Nevertheless, due to CAZJ0211+1051 reaching a historically high radio flux at the time IC131014A arrived, fulfilling the BB95 at the peak of a BBHOP metric in the optical band, and being closer to the most likelihood arrival direction of IC131014A ( of 0.59∘ as opposed to 1.04∘).

### v2

@@ 4 Results and discussion
| Blz | Type | # | Var # | BB95 | Prominent BBHOP | BB95 as BBHOP peak |
| (1) | (2) | (3) | (4) | (5) | (6) | (7) |
| RFC† | (all) | 3225 | 1170 | 8 2.26 (=) | 55 6.19 (=) | 3 1.30 (=) |
| FSRQ | 1773 | 716 | 3 0.81 (=) | 32 3.01 (=) | 0 0.00 (=) | |
| BLL | 296 | 212 | 3 1.31 (=) | 18 2.91 (=) | 2 1.30 (=*) | |
| 4LAC | (all) | 3814 | 1917 | 14 3.44 (=) | 85 11.15 (=) | 4 1.37 (=) |
| FSRQ | 792 | 504 | 1 0.11 (=) | 21 2.18 (=) | 0 0.00 (=) | |
@@
We show the results of the spatiotemporal correlation analysis between the CAZ optical light curves and the IceCube neutrinos in Table 2. We list all individual spatiotemporal associations with (i.e., greater than the median weight) in Table 3. The full list of the spatial and spatiotemporal blazar-neutrino associations is given as an electronic table. The associations are visualized in an all-sky map in Fig. 1.
As described in Sect. 2.2, the test is performed on two samples of blazars: RFC† and 4LAC. Since FSRQs and BLLs are thought to be physically distinct (see Sect. 1), several studies investigated their correlation with neutrinos separately (e.g., Aartsen et al. 2017a). As such, in this paper, we additionally tested the spatiotemporal correlation of neutrinos with FSRQs and BLLs separately. This results in a total of six different blazar subsamples being tested, hence, six rows in Table 2. As described in Sects. 2.2.4, 2.2.5, and 2.2.6, the temporal association is tested via three metrics: (1) BB95 periods; (2) prominent BBHOP flares; and (3) BB95 periods forming the peak of a prominent BBHOP flare. Thus, altogether we have 18 -values.
Overall, we find that the total number of high-weight blazar-neutrino spatiotemporal associations is low. As a result, blazars and neutrinos exhibit a weaker than 2 (pre-trial) spatiotemporal correlation in 14 out of the 18 tested scenarios. Of the four scenarios crossing the 2 threshold, three are weaker than 3, while one reaches 3.48 (pre-trial). Crucially, the strongest correlations are driven by only two high-weight individual associations, namely: (1) the IBL CAZJ0211+1051 (4FGL J0211.2+1051) with the neutrino IC131014A (=); and (2) the IBL CAZJ0509+0541 (TXS 0506+056) with the well-known IC170922A (=).
@@ 4.1 Effect of seasonal gaps
The low number of spatiotemporal associations could be due to a major data limitation present in the CAZ light curves: frequent and long seasonal gaps. Of the 605 spatial associations between neutrinos and variable blazars, in 333 (55.0%) the neutrino arrives in a part of the light curve which is without data. Either the neutrino arrives outside of the total range of the light curve (in 79 cases) or within a seasonal gap (in 172 cases in the CRTS portion, and 82 in the rest of the light curve).
While it is not known which of these spatially neutrino-associated blazars would have been flaring at the time of the neutrino arrival, we can estimate the fraction. For the 272 (45.0% of 605) spatial associations for which we have data available, in 110 (39.7% of 272) of them, the neutrino coincides with a BB95 period or a prominent BBHOP flare (8 with BB95 only, 92 with BBHOP only, and 10 with both BB95 and BBHOP simultaneously). Unsurprisingly, this coincidence rate is similar to the combined average duty cycle of prominent BBHOP flares (34%; see Fig. 12 of K26) and that of BB95 (5%; see Fig. 14 of of K26). If we assume such a flare coincidence rate for all variable light curves globally, we estimate that around potential spatiotemporal associations are missed because of the seasonal gaps. This is a notable caveat for our spatiotemporal analysis because it prevents us from identifying a potential temporal association in around of the spatial associations.
@@ 4.2 BB95 periods versus prominent BBHOP flares
As seen from Table 2, using BB95 periods as a proxy for temporal associations leads to substantially stronger, albeit fewer, correlations than using prominent BBHOP flares. As described in Sect. 2.2.4, BB95 periods are exclusive to the top 5% flux densities and can be as short as a few days. On the other hand, as described in Sect. 2.2.5, BBHOP flares trace a major outburst, whose peak reaches the 75th percentile flux density, from its inception to its termination regardless of the flux density at the start and end. BBHOP flares can last as long as a few hundred days at a time.
@@
Interestingly, the strongest correlation strengths are found upon minimizing the effect of the noise-driven variations. This is done by temporally associating neutrinos with BB95 periods which form the peak of a prominent BBHOP flare (see Sect. 2.2.6). Since this metric requires a relatively high cadence to be identified, they are generally rare in the CAZ light curves. Still, two of the five blazar-neutrino associations emerging from this metric are with the two most high-weight spatiotemporal associations, which greatly enhances the correlation strengths obtained via this temporal metric. This reinforces the above interpretation that neutrinos preferably arrive at the peak of the largest outburst periods from blazars.
@@ 4.3 FSRQs versus BLLs
When testing blazar-neutrino correlations with blazars divided into FSRQs and BLLs, Table 2 shows that BLLs are more strongly spatiotemporally correlated with the updated IceCat1+ neutrinos than FSRQs. In both RFC† and 4LAC samples, and via all the temporal metrics, the correlation strength substantially increases when going from all blazars to BLLs, while a decrease is seen when going from all blazars to FSRQs. This generally contrasts the results that suggest FSRQs are more strongly correlated with neutrinos (e.g., Moretti & Caccianiga 2025).
Naively, our result would suggest that BLLs are better neutrino emitters than FSRQs. However, this generalization should be considered cautiously, because the correlation is driven by only two individual high-weight associations, namely: (1) the IBL CAZJ0211+1051 with IC131014A; and (2) the IBL CAZJ0509+0541 with IC170922A. First, it is possible that we are missing other potential spatiotemporal associations due to the presence of long and frequent seasonal gaps in the CAZ light curves (see Sect. 4.1). With more data, it is possible that this apparent BLL exclusivity disappears. Second, these correlations could arise because of some specific properties present in these individual sources, possibly unique to them rather than their entire class. Third, this generalization is strongly dependent on the reliability of the source classifications. If these blazars are misidentified, they can falsely favor the misidentified class as the more likely neutrino emitter because of their high weight. In this regard, the BLL-classification of some neutrino-emitting candidate blazars have been questioned in the past (e.g., that of TXS 0506+056, Padovani et al. 2019).
@@ 4.4 Fraction of cosmic neutrinos emitted by blazars
In K25 we found that even if only 20% of neutrinos were emitted by blazars at the time of a major flare, our most optimal spatiotemporal test setup (counted and top-hat weighted; used in this study) should confidently detect the correlation. The simulations had 4000 blazars with a 30% flaring rate, comparable to the RFC† and 4LAC tests here. Of the 1000 simulations we ran, in 87.9% the blazar-neutrino spatiotemporal correlation was detected at least at the 4 level, and in 97.8% at the 3 level. The mean of the 1000 -values was 0.0004. To compare the observed significances to the simulation ones, we focus on -values obtained via the BB95 temporal metric. This is because the BB95 temporal metric uses the extreme tail of the flux density distribution to identify flares, similar to how flares were defined in the simulations. Thus, the observed -value for all RFC† blazars is 0.1246 (i.e., 1.54) and for all 4LAC blazars 0.2038 (i.e., 1.27). Clearly, fewer than 20% of the updated IceCat1+ cosmic neutrinos were emitted during major optical flares of blazars.
@@
There are two differences between the tests here and those of K25: (1) the effect of seasonal gaps was not simulated in K25, whereas here the light curves have a seasonal gap rate of 55% leading to the omission of 22% of spatial associations as potential spatiotemporal ones (see Sect. 4.1); and (2) in K25 the original IceCat1+ neutrino sample (283 events) was used, but here we use the updated version (356 events; larger by 26%). While the effect of (1) may be critical and needs to be accounted for, we checked that the effect of (2) is minimal by repeating the simulations of K25 using the updated IceCat1+.
To roughly estimate the fraction of the updated IceCat1+ cosmic neutrinos emitted by blazars during major optical flares, here we run similar simulations to those of K25 while taking into account the seasonal gaps. Within a simulation run, we randomly generate one set of blazars with a sample size of 4000 and a variability rate of 30%, which are generally comparable to those of the 4LAC and RFC† samples. The random generation process is described in detail in Sect. 3.1.1 of K25. In short, we generate 4000 random coordinates, assign each a random following the observed distribution, and finally assign each a random flux density distribution similar to those observed, which is then used to randomly generate flux densities with respect to all neutrinos in the updated IceCat1+. At the end, some of these 4000 random blazars can be spatiotemporally associated with a neutrino by chance, if they happen to fall within the error region of the neutrino, have a critically large , and have a critically large flux density associated with the neutrino. In addition to these, we simulate a number of blazars to be spatiotemporally associated with neutrinos from the updated IceCat1+ (selected randomly based on their signalness). By design, these blazars are randomly centered on the most likely arrival direction of their respective neutrino event with a Gaussian deviation equivalent to the neutrino event error bars, have a critically large , and a critically large flux density associated with the neutrino (see Sect. 3.1.2 of K25). Since the exact number of spatiotemporally neutrino-associated blazars depends on the signalness of the neutrinos, it can vary from one simulation run to another. This number also depends on the fraction of cosmic neutrinos assumed to be coming from blazar flares. In the “realistic” scenario of K25, we assumed a 20% fraction. However, here, we gradually increase this fraction from 0% to 20%.
@@
After each simulation run, we estimate the spatiotemporal correlation strength between the generated sample of blazars and the updated IceCat1+. We focus on the counted, top-hat-weighted scenario, as used in the real tests of this study. We simulate the effect of the 55% light curve gap fraction by randomly excluding 50% of spatiotemporal associations from the -value calculations. Next, we repeat this while incrementally increasing the fraction of blazar-flare-emitted cosmic neutrinos from the updated IceCat1+ sample by 1% from 0% to 20%. These simulations are repeated 500 times for each incremental fraction. We note that the Monte Carlo number for each of the 500 repetitions is , meaning that the smallest obtainable -value is 0.0001 (i.e., 3.9; similar to K25).
@@
In Fig. 2, we plot the correlation strength of the simulations in Gaussian against the fraction of the updated IceCat1+ cosmic neutrinos emitted during major blazar flares in the optical band. Notably, when this fraction is 20%, the typical correlation strengths are 4, similar to those from K25. This suggests that excluding even as much as 50% of the spatiotemporal associations from the -value calculations does not seem to affect the -values by much. Therefore, while the presence of seasonal gaps can substantially reduce the total number of spatiotemporal associations, it only modestly weakens the overall spatiotemporal correlation strengths. Nevertheless, we caution that this conclusion may only be valid within the assumed 0%–20% fraction of the cosmic neutrinos from blazar flares.
@@
In Fig. 2, where the gap fraction is taken as 50%, we additionally show the observed correlation strengths for all RFC† and 4LAC blazars via the most comparable temporal metric (i.e., BB95). These are 1.54 and 1.27, respectively. As made evident from Fig. 2, 1.54 spatiotemporal correlation strengths are most often obtained when 2–8% of the updated IceCat1+ cosmic neutrinos are emitted during major blazar flares. Based on the 4LAC tests, this fraction would be even smaller. This fraction would also be smaller if the light curve gap fraction were smaller than 50%. For example, for a light curve gap fraction of 20%, the observed fraction of the blazar-flare-emitted cosmic neutrinos would be 5%. Thus, we can confidently conclude that the fraction of the updated IceCat1+ cosmic neutrinos emitted during major optical flares of blazars is 8%. Remarkably, this is well in agreement with the previous estimates for the fraction of neutrinos coming from blazars (e.g., Aartsen et al. 2017b; Huber 2019; Oikonomou 2022). Interestingly, even among other neutrino-emitting source candidates, like Seyfert-II AGN, neutrino emission appears to be dominated by only a few sources rather than the whole population (e.g., Saurenhaus et al. 2026).
@@
If indeed 8% of the updated IceCat1+ cosmic neutrinos arise during blazar optical flares, it is then expected that only 12 of the neutrinos were emitted during bright optical flares. Due to the previously discussed light curve limitations, we unfortunately cannot confidently identify most of their potential emitters (apart from CAZJ0211+1051 and CAZJ0509+0541).
4.5 Individual spatiotemporal blazar-neutrino associations
@@
In this subsection, we individually investigate a few interesting spatiotemporal blazar-neutrino associations. We begin with all four associations arising from the BB95 metric with neutrinos that have . This threshold selects the most reliably reconstructed neutrino events and is similar to a common quality threshold used in the literature; namely, and . In the updated IceCat1+ sample there are 79 (22% of 356) neutrino events that meet , of which 64 (81%) are classified as “gold” events by the IceCube collaboration. Their minimum, median, and maximum (and ) are 0.502, 0.663, and 0.997, respectively. Likewise, the minimum, median, and maximum areas () of the enlarged error regions are 3.3, 5.2, and 12.3 , respectively.
@@
Two of the four BB95 associations with (namely, the IBL CAZJ0211+1051 with IC131014A, and the IBL CAZJ0509+0541 with IC170922A) also fulfill the BB95 as the peak of a BBHOP temporal metric. These two associations are the most prominent in our study. In Appendix A, we also explore three additional associations that arise from this metric regardless of their low . The complete list of spatial and spatiotemporal associations is given in an electronic table.
@@
Dedicated studies on CAZJ0211+1051 are needed to confirm if it really is a neutrino emitter. Recent studies have suggested a link between neutrino emission and changes in the multiwavelength polarization degree and angle of blazars (directly, e.g., Paraschos et al. 2025; and indirectly, e.g., Novikova et al. 2023; Blinov & Novikova 2025). Additionally, polarization measurements can reliably trace shocked regions in the jet (e.g., Liodakis et al. 2022b; Kouch et al. 2024b) that could accelerate protons (a key ingredient of unraveling the blazar-neutrino mystery; e.g., Stathopoulos & Petropoulou 2026). Interestingly, CAZJ0211+1051 is one of the prime targets of multiwavelength photo-polarimetric studies aimed at distinguishing between leptonic and hadronic emissions in blazar jets (e.g., Peirson et al. 2022; Kouch et al. 2025c; Agudo et al. 2025) and is known to exhibit extreme, rapid changes in its polarization properties (e.g., Liodakis et al. 2024). Thus, by combining such photo-polarimetric studies with dedicated polarization monitoring, future studies could confirm the suspected neutrino-emitting nature of CAZJ0211+1051.
The next spatiotemporal association we focus on is that of CAZJ0207+0950 (4FGL J0207.9+0953; a blazar candidate) with IC131014A, which is, as discussed above, also spatiotemporally associated with the neighboring blazar CAZJ0211+1051. As seen in Fig. 4, the neutrino arrived 20 d after the peak of a major flare. Notably, while this meets both the BB95 and BBHOP metrics separately, it does not meet the BB95 at the peak of a BBHOP metrics since the neutrino did not arrive exactly at the peak of the BBHOP flare. Due to the statistical nature of the neutrino data, it is not possible to definitively determine from which (if any) of the two sources this neutrino emerged. Nevertheless, due to CAZJ0211+1051 reaching a historically high radio flux at the time IC131014A arrived, fulfilling the BB95 at the peak of a BBHOP metric in the optical band, and being closer to the most likely arrival direction of IC131014A ( of 0.59∘ as opposed to 1.04∘), it could be favored over CAZJ0207+0950 as the emitter of IC131014A.